OpenStudy (anonymous):

I need to find the partivular solution to the differential equation dy/dx=(1+y)/x Given the initial conditon f(-1)=1

OpenStudy (bahrom7893):

I love diff equations! =) So first of all see if its separable

OpenStudy (bahrom7893):

it is!

OpenStudy (bahrom7893):

So separate it: \[dy/dx = (1+y)/x\]

OpenStudy (bahrom7893):

Sorry, I keep crashing on here but i will always reply: So multiply both sides by dx and divide both sides by 1+y

OpenStudy (bahrom7893):

You will have: \[dy/(1+y) = dx/x\]

OpenStudy (anonymous):

can't you think of it as dy/(1+y) you're in good hands later.

OpenStudy (bahrom7893):

Integrate both sides: Left side will be just Ln|1+y| and the right side will be Ln|x|

OpenStudy (bahrom7893):

So: Ln|1+y| = Ln|x| + C

OpenStudy (bahrom7893):

Raise e to both sides to get rid of Ln: \[e^{Ln|1+y|} = e^{Ln|x|+C}\]

OpenStudy (bahrom7893):

Rewrite and simplify: \[e^{Ln|1+y|} = e^{Ln|x|}*e^C\] \[1+y = Kx\]

OpenStudy (bahrom7893):

(e to some constant is another constant so I just let that other constant be K) Now apply initial conditions: f(-1)=1: 1+1 = -1K; 2 = -K; K = -2

OpenStudy (bahrom7893):

So your answer is: 1+y = -2x, or: y = -2x - 1 <= Final answer

OpenStudy (bahrom7893):

Im taking a differential equations course! IT IS FUN!! P.S.: Please click on become a fan if I helped, I really want to get to the next level!! Thanks =)

OpenStudy (anonymous):

why do you multiply by e^c

OpenStudy (bahrom7893):

there's a property: \[A^{B+C} = A^B * A^C\], so in our case: \[e^{Ln|x|+C} = e^{\ln|x|} * e^C\]

OpenStudy (bahrom7893):

and e to some constant is another constant so I said let that constant be K. Any other questions?

OpenStudy (anonymous):

no and thanks for your help